# Revisiting Semi-Supervised Learning in the Era of Foundation Models

### v1

@@ 3.2 A Diverse and Challenging SSL Benchmark
Retinopathy Recog. Medical $36,825$ $5$ $40,80$
Clevr-C Count Synthetic $56,000$ $8$ $10,20$
@@
To effectively evaluate SSL methods in the VFM era, we propose a new benchmark designed to capture the complexities of real-world applications across different tasks, domains, and dataset sizes. Due to these intricacies, frozen VFMs face much greater challenges in our benchmark, as shown in Table 1, reinforcing the necessity of SSL to unleash the potential of VFMs. Specifically, we select six datasets from VTAB (Zhai et al., 2020) covering recognition, depth estimation, and counting tasks, drawn from textual, natural, remote sensing, medical, synthetic, and autonomous driving domains. We also vary label-per-class settings to assess the robustness of SSL methods. We summarize the datasets in Table 2 and provide detailed information inAppendix A. We believe our benchmark provides a more diverse and comprehensive evaluation of SSL methods in the VFM era.
3.3 A Fair Hyperparameter Tuning for SSL
Hyperparameter tuning has long been a persistent challenge in SSL (Oliver et al., 2018), remaining ambiguous and unstandardized throughout much of the existing literature. Due to the labeled data scarcity, standard train-validation split is often infeasible. Tuning on a held-out labeled validation or even test set can lead to significant data leakage, resulting in over-optimistic results and unfair comparisons. To tackle this issue and complement our benchmark, we establish a more rigorous protocol for hyperparameter tuning in SSL.
Our core insight is to harness the defining characteristic of SSL—an abundance of unlabeled training data—to tune parameters in an unsupervised manner and avoid the pitfalls of data leakage. One recent study has explored such an idea for domain adaptation, using unsupervised criteria such as RankMe (Garrido et al., 2023) and AMI (Vinh et al., 2009) to estimate the effectiveness of each hyperparameter configuration (Ericsson et al., 2023).
However, Ericsson et al. (2023) also revealed that no single criterion could reliably select suitable hyperparameters across all scenarios. Motivated by this finding, we propose integrating seven unsupervised criteria—five derived from features, AMI (Vinh et al., 2009), ARI (Halkidi et al., 2002), V-Measure (Rosenberg & Hirschberg, 2007), FMI (Fowlkes & Mallows, 1983) and BNM (Cui et al., 2020), and two from logits, RankMe (Garrido et al., 2023) and CHI (Caliński & Harabasz, 1974)—for more robust tuning. Concretely, for each hyperparameter configuration and its corresponding model, we compute all seven criteria using a held-out unlabeled validation set $\mathcal{D}^{v}\coloneqq\{{\bm{x}}^{v}_{i}\}_{i=1}^{N^{v}}$. Next, we rank every hyperparameter configuration based on each criterion and choose the one achieving the lowest average rank across all. We provide formal definitions and detailed procedures in Section C.1. By eliminating reliance on held-out labeled validation sets, our method mitigates data leakage and promotes a more practical and fair tuning protocol for SSL.
@@ 4 Systematic Evaluation of SSL with VFMs
Evaluation Setup. We focus on two representative VFMs—ViT-B/16 CLIP (Radford et al., 2021) and ViT-B/14 DINOv2 (Oquab et al., 2024)—covering language-image contrastive pre-training and self-supervised pre-training strategies, respectively. We examine four representative SSL methods, including FixMatch (Sohn et al., 2020), FlexMatch (Zhang et al., 2022), SoftMatch (Chen et al., 2023), and FineSSL (Gan & Wei, 2024) and use labeled-only fine-tuning as the baseline. We employ the AdamW optimizer (Loshchilov & Hutter, 2019) with a batch size of 32 and weight decay $5\times 10^{-4}$ to fine-tune the model for 35 epochs. Learning rates and other hyperparameters are tuned with our proposed tuning protocol. The classification performance is evaluated using the Top-1 accuracy on the test dataset.
Figure 2: Average accuracy of SSL methods with full fine-tuning or PEFT (LoRA, AdaptFormer) across 12 settings in our benchmark. Under fair hyperparameter tuning, full fine-tuning with limited labeled data alone can surpass SSL methods. Although PEFT enhances SSL, PEFT with labeled data alone remains highly competitive, indicating limited benefit from unlabeled data in current SSL approaches when paired with VFMs. We present detailed results in Section B.3.
Surprising Effectiveness of Labeled-only Fine-tuning. Surprisingly, under fair comparison, full fine-tuning with even a few labeled images per class can match or surpass SSL methods, as illustrated in Figure 2. In other words, even with a large amount of additional unlabeled data, SSL provides little advantage over fine-tuning VFMs with limited labeled data. This finding is reminiscent of observations made in a comprehensive investigation conducted seven years ago (Oliver et al., 2018).
@@
Can PEFT Come to the Rescue? When labeled downstream data are scarce, recent research has shown that parameter-efficient fine-tuning (PEFT)—which updates only a small subset of parameters or introduces a lightweight learnable module to frozen VFMs—often outperforms full fine-tuning on VFMs (Mai et al., 2024; Xin et al., 2024; Hu et al., 2021; Chen et al., 2022; Jia et al., 2022a; Tu et al., 2023). When considering SSL scenarios with limited labeled data, a natural question arises: is PEFT compatible with SSL? Although a recent study has made an initial attempt to apply PEFT in SSL (Gan & Wei, 2024), a comprehensive analysis of its compatibility with different VFM backbones, PEFT methods, and SSL approaches is still lacking. We, therefore, assess two commonly used PEFT strategies: LoRA (Hu et al., 2021), which trains additive low-rank matrices for Transformer layers to approximate weight updates, and AdaptFormer (Chen et al., 2022), an adapter-based approach proven effective in computer vision. We follow the same protocol outlined in Section 3 to tune the hyperparameters of these PEFT methods.
@@ 5 A Simple, Effective SSL Baseline for VFMs
To wrap up the discussion in Section 4, we propose a new semi-supervised learning method tailored to vision foundation models. Specifically, we build on self-training (Triguero et al., 2015; Lee et al., 2013; Blum & Mitchell, 1998; Balcan et al., 2005; McClosky et al., 2006b, a), a conceptually simple SSL baseline that uses pseudo-labels from unlabeled data as additional supervision to improve the model. In the following section, we first briefly review self-training.
@@ 5.1 Self-training
Let us denote the predicted posterior probability of class $c$ given ${\bm{x}}$ as $p(c|{\bm{x}};f_{\bm{\theta}})$. The core idea of self-training is to gradually assign pseudo-labels to unlabeled data $\mathcal{D}^{u}=\{{\bm{x}}^{u}_{i}\}_{i=1}^{N^{u}}$ based on high-confidence predictions
$\hat{y}^{u}_{i}\coloneqq\operatorname{arg\,max}_{c}p(c|{\bm{x}}^{u}_{i};f_{\bm{\theta}})\hskip 20.0pt\text{ if }\max_{c}p(c|{\bm{x}}^{u}_{i};f_{\bm{\theta}})\geq\tau,$
and add these pseudo-labeled data to the labeled set $\mathcal{D}^{l}=\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{N^{l}}$ for supervised learning. Here, $\bm{\theta}$ represents the weights of the current classifier, and $\tau$ is the confidence threshold. If (a) high-confidence predictions are correct and (b) the updated model gradually increases its confidence in the remaining data, self-training can be as effective as fully supervised learning on the entire dataset with true labels.
In our context, a VFM fine-tuned with PEFT using labeled $\mathcal{D}^{l}$ serves as the current model to generate pseudo-labels for $\mathcal{D}^{u}$, which are then used to further fine-tune the VFM.
Challenges. While self-training is conceptually simple, its practical implementation remains challenging, primarily due to the selection of $\tau$. Ideally, $\tau$ should exclude samples with incorrect pseudo-labels. However, if set too high, it limits the number of pseudo-labeled samples available for supervised learning, reducing effectiveness. Conversely, if $\tau$ is too low, erroneous pseudo-labels may be included, reinforcing model biases. Further complicating the issue, the optimal $\tau$ can vary across self-training iterations. While extensive research has been conducted on pseudo-label selection (Amini et al., 2025; Yang et al., 2022), existing methods are often heuristic or overly complex, inadvertently making self-training impractical or unreliable.
5.2 Our objective
To leverage the high-quality pseudo-labels from label-only PEFT while circumventing the aforementioned challenges, we aim to develop a self-training-based SSL algorithm that is both simple and effective in practice. Specifically, we seek to eliminate the need for complex pseudo-label selection by utilizing all available pseudo-labels (i.e., setting $\tau=0$). This also allows self-training to be completed in a single round, thereby removing an additional hyperparameter.
@@ 5.3 Ensembling of multiple PEFT and VFMs
A necessary step toward achieving our objective is to ensure pseudo-labels of even higher quality to prevent error propagation. To this end, we employ ensembling techniques (Zhou, 2012; Dietterich, 2000), which combine the predictions from multiple models to enhance the overall quality and reliability of pseudo-labels.
@@
What to Ensemble? Unlike conventional methods that train multiple equally performant yet diverse base learners through bootstrapping, random initialization, or cyclic learning rate schedules (Huang et al., 2017), our approach exploits the unique properties of PEFT (Mai et al., 2024) and VFMs (Tong et al., 2024b, a). Specifically, it has been observed that different PEFT methods, while achieving similar overall accuracy on downstream tasks, generate diverse predictions for individual samples. Similarly, VFMs trained on different datasets with varying objective functions exhibit diverse capabilities, with no single model demonstrating clear overall dominance.
How to Ensemble? Given $K$ models $f_{\bm{\theta}_{1}},\dots,f_{\bm{\theta}_{K}}$, several strategies can be applied to integrate their predictions for an unlabeled image ${\bm{x}}^{u}_{i}$. Two common approaches are: (a) Mean Logits, where the logits for each class are averaged across the models; and (b) Mean Probabilities, where the predicted probabilities for each class are averaged. The assumption behind these methods is that the models produce logits or probabilities at similar scales, which is typically the case when the models share similar architectures and are trained on bootstrapped data.
In our context, since different VFMs were trained in distinct ways and adjusted using various PEFT methods, we find that the assumption of similar prediction scales often does not hold, as shown in Figure 3. As a result, some models may dominate the ensembling process, reducing its effectiveness. To address this, we introduce a simple solution called Mean Labels. In this approach, we first obtain “hard” pseudo-labels from each model $\hat{y}_{i,k}^{u}=\operatorname{arg\,max}_{c}p(c|{\bm{x}}_{i}^{u};f_{\bm{\theta}_{k}})$, convert them into one-hot encoding, and average them to obtain a “soft” pseudo-label $\bar{{\bm{e}}}_{i}$. This ensures uniform prediction scales across all models (since the result is always one-hot) while preserving each model’s prediction.
Figure 3: The average entropy of predicted probability distributions from different fine-tuned VFMs. This figure illustrates the gap in entropy among the pseudo-labels generated by different models, indicating that they are not well-calibrated.
5.4 Overall Algorithm (Figure 1)
We summarize our self-training-based SSL algorithm enhanced by ensembling multiple VFMs and PEFT as follows.
Stage 1. Let $n$ denote a specific VFM and $m$ denote a specific PEFT, we fine-tune $\bm{\theta}_{n}$ on $\mathcal{D}^{l}$ to obtain $\bm{\theta}_{n,m}$.
Stage 2. For an unlabeled image ${\bm{x}}_{i}^{u}\in\mathcal{D}^{u}$, we apply $\bm{\theta}_{n,m}$ to obtain its one-hot prediction ${\bm{e}}_{\hat{y}_{i,n,m}^{u}}\in\Delta^{C-1}$, where $\hat{y}_{i,n,m}^{u}=\operatorname{arg\,max}_{c}p(c|{\bm{x}}_{i}^{u};f_{\bm{\theta}_{n,m}})$.
Stage 3. We average the one-hot encoding across all fine-tuned models to obtain the soft pseudo-label
$\bar{{\bm{e}}}_{i}\coloneqq\cfrac{1}{N\times M}\sum_{n}\sum_{m}{\bm{e}}_{\hat{y}_{i,n,m}^{u}}\in\Delta^{C-1}.$ (1)
Stage 4. Let $\mathcal{D}^{p}=\{({\bm{x}}^{u}_{i},\bar{{\bm{e}}}_{i})\}_{i=1}^{N^{u}}$ denote the pseudo-labeled dataset, we re-fine tune a targeted VFM $n\star$ by a targeted PEFT $m\star$ on $\mathcal{D}^{p}$, using the Kullback-Leibler divergence loss function. The resulting $\bm{\theta}_{n\star,m\star}$ is the final SSL model.
@@ 6 Experiment
This section aims to answer the following questions: (1) How does the proposed method perform compared with the existing SSL methods? (2) Does the proposed method scale with more PEFT methods and VFMs? (3) How does the proposed ensemble method perform? (4) How effective is our proposed hyperparameter tuning protocol? We follow the setup described in Section 3 and Section 4.
@@
Figure 4: Incorporating more diverse pseudo-labels for ensemble, ST $\to$ ST(PEFT) $\to$ ST(PEFT+VFMs), boosts performance, highlighting the benefits of pseudo-label diversity.
6.1 Performance Comparison
Following Section 5.4, we use ST(PEFT+VFMs) to refer to ensembling over multiple PEFT methods and VFMs; ST(PEFT) to refer to ensembling over multiple PEFT methods but one VFM; and ST to refer to no ensembling. In all cases, when we fine-tune VFMs using pseudo-labels, we re-initialize the model from the original VFMs.
Section 6 shows the performance comparison between ST, ST(PEFT), ST(PEFT+VFMs) and existing SSL methods across all the 12 settings in our benchmark. Overall, ST(PEFT+VFMs) achieve superior performance compared with others, as shown in the Average column. To further analyze the results, Figure 5 visualize the ranking frequency, where each matrix entry $(i,j)$ indicates how often method $i$ is ranked $j$-th across 12 settings. We then compute the mean rank (number in the bracket) of each method and sort them accordingly. Although ST(PEFT+VFMs) does not claim first place in every single setting, it secures the top rank most frequently, showcasing the effectiveness of pseudo-label ensemble and establishing it as a simple yet powerful baseline in the VFM era.
Digging deeper into the three ST-based methods, we observe a consistent performance boost as more diverse pseudo-labels are introduced for the ensemble, i.e., ST $\to$ ST(PEFT) $\to$ ST(PEFT+VFMs), as shown in Figure 4. This underscores the importance of diversity among pseudo-labels. An open question emerging from these results is whether the benefits will continue to scale as even more PEFT methods and VFMs are integrated.
Figure 5: Ranking frequency of different SSL methods in the proposed benchmark. The number in $(i,j)$ indicates the frequency of method $i$ is ranked $j$-th across 12 settings. The number in the bracket indicates the average rank.

### v2

@@ 3.3 Fair Hyperparameter Tuning for SSL
Retinopathy Recog. Medical 36,825 5 4, 8
Clevr-C Count Synthetic 56,000 8 1, 2
@@
Our core insight is to harness the defining characteristic of SSL—an abundance of unlabeled training data—to tune parameters in an unsupervised manner and avoid the pitfalls of data leakage. One recent study has explored such an idea for domain adaptation, using unsupervised criteria such as RankMe garrido2023rankmeassessingdownstreamperformance and AMI vinh2009information to estimate the effectiveness of each hyperparameter configuration ericsson2023betterpracticesdomainadaptation .
However, ericsson2023betterpracticesdomainadaptation also revealed that no single criterion could reliably select suitable hyperparameters across all scenarios. Motivated by this finding, we propose integrating seven unsupervised criteria—five derived from features, AMI vinh2009information , ARI halkidi2002cluster , V-Measure rosenberg2007v , FMI fowlkes1983method and BNM cui2020towards , and two from logits, RankMe garrido2023rankmeassessingdownstreamperformance and CHI calinski1974dendrite —for more robust tuning. Concretely, for each hyperparameter configuration and its corresponding model, we compute all seven criteria using a held-out unlabeled validation set $\mathcal{V}\coloneqq\{{\bm{x}}^{v}_{i}\}_{i=1}^{|\mathcal{V}|}$. Next, we rank every hyperparameter configuration based on each criterion and choose the one achieving the lowest average rank across all. We provide formal definitions and detailed procedures in section C.1 and the effectiveness of the proposed method in section B.7. By eliminating reliance on held-out labeled validation sets, our method mitigates data leakage and promotes a more practical and fair tuning protocol for SSL.
@@ 4 Systematic Evaluation of SSL with VFMs
Evaluation Setup. We focus on two representative VFMs—ViT-B/16 CLIP radford2021learningtransferablevisualmodels and ViT-B/14 DINOv2 oquab2024dinov2learningrobustvisual —covering language-image contrastive pre-training and self-supervised pre-training strategies, respectively. We examine four representative SSL methods, including FixMatch sohn2020fixmatch , FlexMatch zhang2021flexmatch , SoftMatch chen2023softmatch , and FineSSL gan2024erasingbiasfinetuningfoundation and use labeled-only fine-tuning as the baseline. We employ the AdamW optimizer Loshchilov2019Decoupled with a batch size of 32 and weight decay $5\times 10^{-4}$ to fine-tune the model for 35 epochs. Learning rates and other hyperparameters are tuned with our proposed tuning protocol. The classification performance is evaluated using the Top-1 accuracy on the test dataset.
Surprising Effectiveness of Labeled-only Fine-tuning. Surprisingly, under fair comparison, full fine-tuning with even a few labeled images per class can match or surpass SSL methods, as illustrated in Figure 4. In other words, even with a large amount of additional unlabeled data, SSL provides little advantage over fine-tuning VFMs with limited labeled data. This finding is reminiscent of observations made in a comprehensive investigation conducted seven years ago oliver2018realistic .
@@
Can PEFT Come to the Rescue? When labeled downstream data are scarce, recent research has shown that parameter-efficient fine-tuning (PEFT)—which updates only a small subset of parameters or introduces a lightweight learnable module to frozen VFMs—often outperforms full fine-tuning on VFMs mai2024lessonslearnedunifyingempirical ; vpetlbench2024 ; hu2021loralowrankadaptationlarge ; chen2022adaptformeradaptingvisiontransformers ; jia2022visual ; tu2023visual . When considering SSL scenarios with limited labeled data, a natural question arises: is PEFT compatible with SSL? Although a recent study has made an initial attempt to apply PEFT in SSL gan2024erasingbiasfinetuningfoundation , a comprehensive analysis of its compatibility with different VFM backbones, PEFT methods, and SSL approaches is still lacking. We, therefore, assess two commonly used PEFT strategies: LoRA hu2021loralowrankadaptationlarge , which trains additive low-rank matrices for Transformer layers to approximate weight updates, and AdaptFormer chen2022adaptformeradaptingvisiontransformers , an adapter-based approach proven effective in computer vision. We follow the same protocol outlined in section 3 to tune the hyperparameters of these PEFT methods.
@@ 5 A Simple, Effective SSL Baseline for VFMs
To wrap up the discussion in section 4, we propose a new semi-supervised learning method tailored to vision foundation models. Specifically, we build on self-training triguero2015self ; lee2013pseudo ; blum1998combining ; balcan2005co ; mcclosky2006reranking ; mcclosky2006effective , a conceptually simple SSL baseline that uses pseudo-labels from unlabeled data as additional supervision to improve the model. In the following section, we first briefly review self-training.
@@ 5.1 Self-training
Let us denote the predicted posterior probability of class $c$ given ${\bm{x}}$ as $p(c\big{|}{\bm{x}};f({\bm{\theta}}))$. The core idea of self-training is to gradually assign pseudo-labels to unlabeled data $\mathcal{U}=\{{\bm{x}}^{u}_{i}\}_{i=1}^{|\mathcal{U}|}$ based on high-confidence predictions
[…]
and add these pseudo-labeled data to the labeled set =$\mathcal{L}=\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{|\mathcal{L}|}$ for supervised learning. Here, $\bm{\theta}$ represents the weights of the current classifier, and $\tau$ is the confidence threshold. If (a) high-confidence predictions are correct and (b) the updated model gradually increases its confidence in the remaining data, self-training can be as effective as fully supervised learning on the entire dataset with true labels.
In our context, a VFM fine-tuned with PEFT using labeled $\mathcal{L}$ serves as the current model to generate pseudo-labels for $\mathcal{U}$, which are then used to further fine-tune the VFM.
Challenges. Although self-training is conceptually simple, choosing the confidence threshold $\tau$ is difficult. A high $\tau$ excludes wrongly labeled samples but leaves too few for learning; a low $\tau$ admits errors that reinforce bias. Moreover, the optimal $\tau$ shifts across iterations. Despite extensive study of pseudo-label selection amini2025self ; yang2022survey , most approaches remain heuristic or overly complex, undermining self-training’s practicality and reliability.
Our objective. To leverage the high-quality pseudo-labels from label-only PEFT while circumventing the aforementioned challenges, we aim to develop a self-training-based SSL algorithm that is both simple and effective in practice. Specifically, we seek to eliminate the need for complex pseudo-label selection by utilizing all available pseudo-labels (i.e., setting $\tau=0$). This also allows self-training to be completed in a single round, thereby removing an additional hyperparameter.
@@ 5.2 Ensembling of multiple PEFT and VFMs
Input :
Labeled dataset $\mathcal{L}$ and unlabeled dataset $\mathcal{U}$
PEFT methods indexed by $n\in[1,N]$
VFMs indexed by $m\in[1,M]$
Initialized parameters of VFM on PEFT ${\bm{\theta}}_{n,m}$
Output : Optimal parameter $\bm{\theta}^{*}$
// (a) Supervised Parameter Efficient Fine-Tuning
for $n\in[1,N],\,m\in[1,M]$ do
       Fine-tune ${\bm{\theta}}_{n,m}$ on $\mathcal{L}$ to obtain $\tilde{\bm{\theta}}_{n,m}$
// (b) Pseudo-Label Generation
for $n\in[1,N],\,m\in[1,M]$ do
       Compute one-hot pseudo-label set $\mathcal{P}_{n,m}$ as
[…]
// (c) Pseudo-Label Ensemble
Ensemble pseudo-label $\mathcal{P}\coloneqq\{\bar{p}_{i}\}_{i}^{|\mathcal{U}|}$, where:
$\bar{p}_{i}=\frac{1}{N\times M}\sum_{\bar{n}\in[1,N],\bar{m}\in[1,M]}\mathcal{P}_{\bar{n},\bar{m}}[i]$
// (d) Self-Training
Choose $n^{\star}\in[1,N],\,m^{\star}\in[1,M]$ and fine-tune ${\bm{\theta}}_{n\star,m\star}$ on $\mathcal{P}$ to get $\bm{\theta}^{*}$
Algorithm 1 V-PET (Figure 2)
@@
A necessary step toward achieving our objective is to ensure pseudo-labels of even higher quality to prevent error propagation. To this end, we employ ensembling techniques zhou2012ensemble ; dietterich2000ensemble , which combine the predictions from multiple models to enhance the overall quality and reliability of pseudo-labels.
@@
What to Ensemble? Unlike conventional methods that train multiple equally performant yet diverse base learners through bootstrapping, random initialization, or cyclic learning rate schedules huang2017snapshot , our approach exploits the unique properties of PEFT mai2024lessonslearnedunifyingempirical and VFMs tong2024eyes ; tongcambrian . Specifically, it has been observed that different PEFT methods, while achieving similar overall accuracy on downstream tasks, generate diverse predictions for individual samples. Similarly, VFMs trained on different datasets with varying objective functions exhibit diverse capabilities, with no single model demonstrating clear overall dominance.
How to Ensemble? Common strategies for ensembling different predictions are (a) Mean Logits—averaging class logits, and (b) Mean Probabilities—averaging class probabilities, both assuming similar output scales (typical with identical architectures and bootstrapped training). However, VFMs trained differently and fine-tuned via varied PEFT methods produce inconsistent scales (Figure 4), causing some models to dominate and weakening the ensemble.
To address this, we introduce a simple solution called Mean Labels. We first obtain predictions from each model, convert them into one-hot encoding to ensure uniformity, and average them to obtain a soft pseudo-label. Finally, we generate the augmented dataset $\mathcal{P}$ using averaged soft pseudo-label and fine-tune PEFT-based VFM models. We present our proposed pipeline in algorithm 1.
@@ 6 Experiment
We conduct experiments to address four questions—(1) How does our method compare to existing SSL? (2) Does it scale with more PEFTs and VFMs? (3) How effective is the ensemble strategy? Following the setups in section 3 and section 4, we evaluate two variants: V-PET (ensembles over PEFT methods and VFMs) and PET (ensembles over PEFT methods on a single VFM). We use ST (self-training without ensembling) as a baseline and, in all cases, reinitialize from the original pretrained VFM before fine-tuning with pseudo-labels.
Figure 3: Average SSL accuracy with full fine-tuning or PEFT across 12 settings shows that, with fair hyper-parameter tuning, fine-tuning on limited labels can outperform SSL; PEFT boosts SSL yet matches labeled-only performance, indicating minimal unlabeled-data benefit in current VFM-based SSL (see section B.3).
Figure 4: The average entropy of predicted probability distributions from different fine-tuned VFMs, highlighting the entropy gap among pseudo-labels, indicating poor calibration.
6.1 Performance Comparison
We summarize the performance comparison between ST, PET, V-PET and existing SSL methods across all the 12 settings in subsection 6.1. Overall, V-PET achieve superior performance compared with others, as shown in the Average column. To further analyze the results, Figure 6 visualize the ranking frequency, where each matrix entry $(i,j)$ indicates how often method $i$ is ranked $j$-th across 12 settings. We then compute the mean rank (number in the bracket) of each method and sort them accordingly. Although V-PET does not claim first place in every single setting, it secures the top rank most frequently, showcasing the effectiveness of pseudo-label ensemble and establishing it as a simple yet powerful baseline in the VFM era.
@@ 6.1 Performance Comparison
Digging deeper into the three ST-based methods, we observe a consistent performance boost as more diverse pseudo-labels are introduced for the ensemble, i.e., ST $\to$ PET $\to$ V-PET, as shown in Figure 1. This underscores the importance of diversity among pseudo-labels.